Nachman Industries just paid a dividend of D0 = $3.75. Analysts expect the company's dividend to...

  1. Nachman Industries just paid a dividend of D0 = $3.75. Analysts expect the company's dividend to grow by 30% this year, by 10% in Year 2, and at a constant rate of 5% in Year 3 and thereafter. The required return on this stock is 9.00%. What is the best estimate of the stock’s current market value?
Answer #1

D1=(3.75*1.3)=4.875

D2=(4.875*1.1)=5.3625

Value after year 2=(D2*Growth Rate)/(Required return-Growth Rate)

=(5.3625*1.05)/(0.09-0.05)

=$140.765625

Hence current value=Future dividends and value*Present value of discounting factor(rate%,time period)

=4.875/1.09+5.3625/1.09^2+$140.765625/1.09^2

which is equal to

=$127.47(Approx).